OpenStudy (anonymous):

find the fist derivative to f(x)=1+2e^-x+xe^-x

OpenStudy (anonymous):

\[f(x)=1+2e ^{-x}-xe ^{-x}\]

OpenStudy (anonymous):

is it the first one or the second one?

OpenStudy (anonymous):

the first

OpenStudy (shamil98):

\[f(x) = 1 + 2e^{-x} - xe^{-x}\] \[f'(x) = -2e^{-x} -e^{-x} + x e^{-x}\]

OpenStudy (anonymous):

Ok! The answer is: \[f'(x)=-2e ^{-x}-xe^{-x}+e^{-x}\]

OpenStudy (shamil98):

you use the product rule for the -xe^{-x}

OpenStudy (anonymous):

also: \[(e^{u})'=u'\times e^{u}\]

OpenStudy (shamil98):

@lego_less check your answer.. it is wrong..

OpenStudy (anonymous):

No ur answer is right.